# ZeroTuning: Unlocking the Initial Token’s Power to Enhance Large Language Models Without Training

### v1

@@ 2.4 Analyzing the Role of the Initial Token Across Attention Heads
Unlike layers, which pass information sequentially, attention heads operate in parallel and contribute independently via concatenation. It remains unclear how they differ in response to the initial token. To investigate this, we increase the initial token’s attention of each head individually by applying $\gamma=1.5$, and evaluate the model’s performance on SST-2 and MMLU. For comparison, we also evaluate (i) no scaling ($\gamma=1$), (ii) uniform up-scaling ($\gamma=1.5$) across all heads, and (iii) uniform down-scaling ($\gamma=0.6$) across all heads.
@@
These findings align with prior work suggesting that attention heads specialize into distinct functional roles during pretraining [27]. Such roles include global retrieval, structural parsing, option discrimination, and sensitivity to negation. We hypothesize that these functional differences drive the varying impact of initial token attention scaling. While some heads rely on broad attention to support global reasoning, others benefit from focused attention on salient tokens. Together, these specialized mechanisms enable language models to better understand and process diverse input patterns.
@@ 2.6 ZeroTuning Methodology
Building on our empirical findings, we propose ZeroTuning, a simple yet effective method that adjusts the initial token’s attention in a head-specific manner to enhance LLM performance, without requiring task-specific token identification. ZeroTuning comprises three steps:
@@
    Head Behavior Profiling: Using a small calibration set (e.g., a few validation examples), we assess each attention head’s sensitivity to scaling the initial token’s attention. A head is classified as up-effective if increased attention improves performance, and down-effective otherwise.
@@
    Selective Rescaling: Based on the dominant head type (i.e., the group with the highest proportion), we apply a scaling factor $\gamma$ exclusively to that group to adjust the initial token’s attention. [^1] We explore the effect of scaling different subsets of heads; results (Figure 8, Appendix F) show that tuning the top 20% to 70% yields better performance.
  3. 3.
    Renormalization: The scaled attention scores are re-normalized using the softmax function to maintain a valid distribution.
For optimized attention implementations (e.g., SDPA, Flash Attention) where direct modification of attention scores is infeasible, ZeroTuning applies scaling to the query or key states before attention computation. We demonstrate that tuning key states yields similar effects to attention-based tuning, but induces steeper changes in the attention distribution of subsequent tokens (see Appendix A). The following sections evaluate the performance and robustness of ZeroTuning across different tasks and model settings.
@@ 3.1 Experimental Setup
Models, Tasks, and Datasets. Models: We evaluate ZeroTuning on four LLMs with distinct attention implementations: Llama-3.1-8B-Instruct [28] and Llama-2-13B-Chat [29] with eager attention, Qwen-2-7B [30] with SDPA attention, and DeepSeek-R1-14B [31] with Flash attention. [^2] Eager, SDPA, and Flash are official attention implementations in modern Transformer libraries. Eager computes the full attention map; SDPA uses PyTorch’s efficient API to select the optimal implementation; Flash relies on fused CUDA kernels from the FlashAttention library. Tasks and Datasets: Our experiments encompass three task types across 15 datasets: (1) Text Classification and Reasoning, including SST-2 (binary sentiment classification) [32], SST-5 (fine-grained sentiment analysis) [32], MR (movie review polarity detection) [33], SUBJ (subjectivity classification) [34], TREC (question type classification) [35], CB (commitment detection) [36], and BoolQ (boolean question answering) [37]; (2) Domain-Specific Multiple-Choice, including MMLU (cross-domain knowledge testing) [38], AQUA (math word problems) [39], MathQA (algebraic reasoning) [40], LogiQA (logical reasoning) [41], CQA (commonsense reasoning) [42], PIQA (physical commonsense QA) [43], and ARCC (scientific reasoning) [44]; and (3) Multi-Round Conversation, using MT-Bench [39].
Baselines and Evaluation Metrics.
Baselines: We benchmark ZeroTuning against three baselines: (1) vanilla inference, which performs standard inference without any modifications; (2) ACT [12], which identifies sink tokens using an attention score threshold and reduces their attention weights; and (3) Auto-PASTA [10], which leverages an LLM to locate important tokens and enhance their attention weights. For ACT, we use the original hyperparameters ($\beta=0.4$) and filter attention heads on a 10% validation set. For AutoPASTA, the evaluated LLM identifies key input regions based on prompts provided in Appendix J. Since ACT relies on explicit attention maps, it is only evaluated on LLaMA-3.1-8B-Instruct with eager attention implementation. Evaluation Metrics: We assess performance using accuracy for text classification and multiple-choice tasks. For the multi-round conversation task, we report average quality scores as evaluated by GPT-4, following the methodology outlined in [39].
Implementation Details. ZeroTuning is implemented using PyTorch and the Hugging Face Transformers library. We select zero-shot setting with standard greedy decoding for baseline consistency, and use prompt templates from [45, 46, 47] across datasets (see Appendix J for details). Unless otherwise specified, we tune the top 40% of the heads. Each experiment uses a randomly selected validation set of 500 samples with the random seed fixed to 42.
Table 1: Performance Comparison of Classification Tasks Across Models. The best performance in each dataset is bolded, the second best is underlined, and the ZeroTuning method is highlighted in gray.
Model Method Datasets SST-2 SST-2 MR BoolQ CB TREC SUBJ Avg. Llama-3.1-8B-Instruct Vanilla 73.20 45.40 89.20 69.60 82.14 14.00 44.60 59.59 ACT 85.00 43.80 90.80 58.60 82.14 15.80 44.60 60.11 Auto-PASTA 89.60 47.20 91.40 72.60 83.93 16.00 45.40 63.73 ZeroTuning 91.60 52.00 92.00 82.40 89.29 26.20 66.60 71.44 Qwen-2-7B (SDPA) Vanilla 78.80 45.40 72.40 85.00 78.50 12.60 13.00 55.10 ACT / / / / / / / / Auto-PASTA 89.00 47.00 77.70 85.00 89.29 14.00 57.00 65.57 ZeroTuning 89.60 47.20 87.40 86.40 85.71 26.60 54.40 68.19 Deepseek-R1-14B (Flash) Vanilla 91.20 49.40 89.20 83.40 89.29 20.80 50.40 67.67 ACT / / / / / / / / Auto-PASTA 92.00 52.20 89.80 83.40 92.86 22.60 50.40 69.04 ZeroTuning 93.00 51.20 90.20 88.00 92.86 32.00 55.80 71.87
Table 2: Performance Comparison of Multiple-Choice Tasks Across Models
@@ 3.2 Overall performance of ZeroTuning
We first evaluate ZeroTuning on various text classification datasets using different LLMs, as shown in Table 1. Despite tuning only a single token, ZeroTuning consistently outperforms baselines and methods that require tuning more tokens. With Llama-3.1-8B-Instruct, it achieves an average improvement of 11.71% over vanilla, with peaks of 22.00% on SUBJ and 18.40% on SST-2. It outperforms AutoPASTA by an average of 7.71%. On Qwen-2-7B, ZeroTuning gains 13.09%, and on Deepseek-R1-14B, it improves by 4.20%, with a notable increase of 11.20% on TREC.
@@
Next, we evaluate ZeroTuning on common domain-specific multiple-choice datasets under various settings, as shown in Table 2. For Llama-3.1-8B-Instruct, it increases the average accuracy by 2.64%, with gains of 3.40% on LogiQA and 1.40% on MMLU. Qwen-2-7B gains 1.74%, and Deepseek-R1-14B gains 2.15%, with an outstanding 7.80% on LogiQA.
@@
We further demonstrate ZeroTuning’s effectiveness in multi-round conversations using MT-Bench [48], with results in Table 3. For Llama-3.1-8B-Instruct, ZeroTuning improves the average score by 0.162 points (7.966 vs. 7.804). For Llama-2-13B-Chat, it achieves a 0.266 points gain (6.916 vs. 6.650), showing its effectiveness in interactive settings
@@
4 Further Analysis
We provide additional analyses to deepen the understanding of ZeroTuning’s behavior and practical utility under varying conditions. (1) Tuning Different Matrices: Appendix A begins with a comparison between tuning attention scores and key states. Tuning attention scores yields stable improvements due to its linear influence, whereas tuning key states is sensitive and can cause sharp performance drops because of exponential effects. (2) Resource Constraints: Appendix B evaluates ZeroTuning across three simulated computational budgets, revealing consistent gains from minimal scaling (Level 0) to full head classification and parameter search (Level 2). (3) Context Length Sensitivity: Appendix C tests robustness to long contexts with irrelevant tokens, showing ZeroTuning maintains accuracy where vanilla models degrade. (4) Few-Shot Learning Robustness: Appendix D demonstrates stable few-shot gains and fewer invalid outputs, indicating ZeroTuning complements in-context learning. (5) Prompt Variation Sensitivity: Appendix G assesses prompt modifications such as dropped or altered instructions; ZeroTuning sustains strong performance. (6) Effect of Attention Head Subset Selection: Appendix F shows that tuning a moderate proportion of attention heads (20%–70%) outperforms both full and minimal tuning. (7) Quantization Robustness: Appendix H shows a consistent trend across 16-, 8-, and 4-bit quantization, with ZeroTuning partially mitigating degradation from low-precision quantization. (8) Decoding Strategy Impact: Appendix E evaluates various decoding methods (Top-$k$, Top-$p$, Beam Search), confirming ZeroTuning’s effectiveness.
5 Related Work
5.1 Token-Level Attention Tuning
Token-level attention tuning typically aims to increase attention to critical input tokens or decrease attention to less informative tokens. [49] proposes a mask perturbation method to adjust attention weights for key tokens, thereby improving translation quality. [11] introduce PASTA, which allows manual designation of important tokens during inference. This is extended by AutoPASTA [10], which uses LLMs to autonomously identify salient tokens and increase attention to them. In contrast, ACT [12] reduces attention to semantically trivial sink tokens and redirects it to meaningful content. Similar strategies have been applied to VLMs to mitigate hallucinations. PAI [13] enhances attention to image tokens at inference time to counteract text-dominant bias. IBD [14] and OPERA [15] further refine this idea by prioritizing visual information or penalizing overconfident summary tokens. While effective, these methods depend on identifying task-specific tokens, which may introduce bias (e.g., overemphasizing misleading tokens) and limit applicability when token importance is unclear or attention maps are unavailable. In contrast, our method focuses on a task-invariant initial token, removing the need for costly token identification, and can be easily applied by tuning key states.
5.2 The Magic of the Initial Token
Recent studies highlight the significance of the initial token, especially through the lens of the attention sink phenomenon, where it draws substantial attention despite low semantic content. [16] show that preserving such tokens is critical for maintaining performance in sliding window attention. [17] attribute this effect to softmax normalization and causal masking, while [18] and [19] identify architectural biases that amplify attention to the initial token, including key-query alignment and LayerNorm effects. Functionally, the initial token is hypothesized to serve as a stabilizing “no-op” anchor, enhancing robustness to prompt variations [19]. It has been leveraged in applications such as long-context modeling [50, 16], but also poses challenges for quantization due to its high attention weight [2, 51]. While previous work has identified the structural and functional importance of the initial token, its potential as a target for attention tuning remains underexplored. In this work, we provide a detailed analysis of attention tuning of the initial token across layers and heads, demonstrating its consistent influence across different tasks. Our approach bridges the gap between the these lines of research by proposing a novel method that advances interpretable attention tuning.

### v2

@@ 3.4 Analyzing the Role of the Initial Token Across Attention Heads
Unlike layers passing information sequentially, attention heads operate in parallel and contribute independently via concatenation. It remains unclear how they differ in response to the initial token. To investigate this, we increase the initial token’s attention of each head individually by applying $\gamma=1.5$, and evaluate the model’s performance on SST-2 and MMLU. For comparison, we also evaluate (i) no scaling ($\gamma=1$), (ii) uniform up-scaling ($\gamma=1.5$) across all heads, and (iii) uniform down-scaling ($\gamma=0.6$) across all heads.
@@
These results align with prior studies showing that attention heads specialize into distinct functional roles during pretraining (Zheng et al., 2024b; Guo et al., 2024), such as global retrieval, structural parsing, option discrimination, and negation sensitivity. We propose that these functional differences may explain the variable impact of initial token attention scaling, with some heads supporting broad global reasoning and others focusing on salient tokens. This interpretation requires further exploration in future work.
@@ 3.6 ZeroTuning Methodology
Building on our empirical findings, we propose ZeroTuning, a method that enhances LLM performance via head-specific attention adjustments to the initial token, without requiring task-specific token identification. The core methodology involves three steps: head behavior profiling, selective rescaling, and renormalization. To accomplish the first two steps, hyperparameter calibration, we introduce both supervised and unsupervised variants to cater to different application scenarios.
Supervised Calibration
Consistent with standard practices in inference-time adaptation (Yu et al., 2024; Zhang et al., 2023a; 2024a), our primary approach utilizes a labeled calibration set (e.g., validation examples) to perform:
@@
    Head Behavior Profiling: Assessing each attention head’s sensitivity to the initial token’s attention scaling. A head is classified as up-effective if increased attention improves accuracy, and down-effective otherwise.
@@
    Selective Rescaling: Applying a scaling factor $\gamma$, identified by searching for the value that maximizes accuracy on the calibration set, exclusively to the dominant head type (i.e., the most numerous group).
Unsupervised Calibration via Entropy Minimization
To mitigate the reliance on labeled data, we propose a novel unsupervised calibration strategy. Based on our finding that a model’s output entropy strongly correlates with its accuracy (as detailed in Section 3.2), this variant bypasses the calibration set entirely. Instead, it identifies the optimal heads and scaling factor $\gamma$ by minimizing the average next-token prediction entropy directly on the unlabeled test inputs.
Finally, the third step, Renormalization, is applied in both variants by re-normalizing the scaled attention scores via the softmax function to maintain a valid probability distribution. For optimized attention implementations (e.g., SDPA, Flash Attention) where direct score modification is infeasible, ZeroTuning applies scaling to the query or key states. We demonstrate and show that this yields similar effects in Appendix G.
@@ 4.1 Experimental Setup
Our evaluation includes four recent LLMs with distinct attention implementations (Llama-3.1-8B, Llama-2-13B, Qwen-2-7B, and DeepSeek-R1-14B). We test performance across 15 datasets spanning three categories: Text Classification, Multiple-Choice QA, and Multi-Round Conversation. We benchmark ZeroTuning against three methods: (1) Vanilla inference; (2) ACT (Yu et al., 2024), a sink-token down-scaling method  [^1] Since ACT explicitly manipulates attention maps, we only evaluate it on Llama-3.1-8B-Instruct; and (3) Auto-PASTA (Zhang et al., 2024a), an LLM-guided key-token up-scaling method. All experiments are conducted in a zero-shot setting with greedy decoding for fair comparison. For our supervised variant and the baselines, hyperparameters are calibrated on a fixed validation set. A detailed description of all models, datasets, baselines, and implementation specifics is provided in Appendix B.
Table 1: Performance Comparison of Classification Tasks Across Models. The best performance in each dataset is bolded and the ZeroTuning method is highlighted in gray.
Model Method Datasets
SST2 SST5 MR BoolQ CB TREC SUBJ Avg.
Llama-3.1-8B-Instruct Vanilla 73.20 45.40 89.20 69.60 82.14 14.00 44.60 59.59
ACT 85.00 43.80 90.80 58.60 82.14 15.80 44.60 60.11
Auto-PASTA 89.60 47.20 91.40 72.60 83.93 16.00 45.40 63.73
[…]
Qwen-2-7B (SDPA) Vanilla 78.80 45.40 72.40 85.00 78.50 12.60 13.00 55.10
ACT
Auto-PASTA 89.00 47.00 77.70 85.00 89.29 14.00 57.00 65.57
[…]
Deepseek-R1-14B (Flash) Vanilla 91.20 49.40 89.20 83.40 89.29 20.80 50.40 67.67
ACT
Auto-PASTA 92.00 52.20 89.80 83.40 92.86 22.60 50.40 69.04
[…]
Table 2: Performance Comparison of Multiple-Choice Tasks Across Models.
@@ 4.2 Overall performance of ZeroTuning
For a fair and direct comparison with existing supervised baselines, our main experiments focus on the supervised ZeroTuning approach.
@@
We first evaluate ZeroTuning on various text classification datasets using different LLMs, as shown in Table 1. Despite tuning only a single token, ZeroTuning consistently outperforms baselines and methods that require tuning more tokens. With Llama-3.1-8B-Instruct, it achieves an average improvement of +11.71% over vanilla, with peaks of +22.00% on SUBJ and +18.40% on SST-2. It outperforms AutoPASTA by an average of 7.71%. On Qwen-2-7B, ZeroTuning gains +13.09%, and on Deepseek-R1-14B, it improves by +4.20%, with a notable increase of +11.20% on TREC.
@@
Next, we evaluate ZeroTuning on common domain-specific multiple-choice datasets under various settings, as shown in Table 2. For Llama-3.1-8B-Instruct, it increases the average accuracy by +2.64%, with gains of +3.40% on LogiQA and +1.40% on MMLU. Qwen-2-7B gains +1.74%, and Deepseek-R1-14B gains +2.15%, with an outstanding +7.80% on LogiQA.
@@
We further demonstrate ZeroTuning’s effectiveness in multi-round conversations using MT-Bench (Zheng et al., 2024a), with results in Table 3. For Llama-3.1-8B-Instruct, ZeroTuning improves the average score by 0.162 points (7.966 vs. 7.804). For Llama-2-13B-Chat, it achieves a 0.266 points gain (6.916 vs. 6.650), showing its effectiveness in interactive settings
@@
4.3 Unsupervised ZeroTuning
We validate the fully unsupervised variant of ZeroTuning, which eliminates the need for any labeled calibration data by minimizing the model’s average output entropy. Appendix E provides a detailed empirical validation for this approach, including visual comparisons, empirical analysis, and a breakdown of error patterns. As shown in Figure 8, this entropy-guided method achieves performance highly competitive with its supervised counterpart, extending its applicability to label-scarce scenarios.
Figure 8: Performance comparison of Vanilla, Supervised, and Unsupervised ZeroTuning on Llama-3.1-8B-Instruct.
5 Further Analysis and Key Findings
Robustness Across Diverse Conditions.
ZeroTuning demonstrates remarkable stability. It maintains strong performance gains even when faced with: (1) long contexts padded with irrelevant distractors, where it stabilizes information flow better than the vanilla model (Appendix I); (2) few-shot scenarios, where it consistently improves instruction-following and reduces invalid outputs (Appendix J); (3) significant prompt variations, including missing or altered instructions (Appendix M); and (4) low-precision 4-bit and 8-bit quantization, where it partially mitigates the associated accuracy degradation (Appendix N).
Practicality and Methodological Choices.
Our method is not only robust but also highly practical. We theoretically and empirically confirm that tuning key states provides a viable, kernel-agnostic alternative to direct attention score manipulation, proving effective in optimized environments like FlashAttention (Appendix G). Furthermore, ZeroTuning is adaptable to resource-constrained settings, delivering gains even with a minimal, search-free scaling approach (Appendix H). We also analyze key methodological choices, showing that tuning a moderate subset of heads (40%–70%) is optimal, providing a clear and efficient configuration (Appendix L).
Boundaries of Efficacy.
Finally, we analyze the method’s boundaries and potential side effects. We quantitatively demonstrate that ZeroTuning excels at correcting a model’s uncertain errors but cannot override high-confidence mistakes rooted in flawed pretrained knowledge. We also characterize the negative effects of extreme tuning, which provides a clearer picture of the method’s operational limits (Appendix D). This positions our method as a powerful tool for unlocking a model’s latent knowledge, rather than a substitute for fine-tuning. Intriguingly, we also find that within a safe operational range, the scaling factor can modulate output diversity in a manner analogous to temperature, but with the unique ability to alter the rank-ordering of logits and thereby correct errors that temperature scaling cannot fix.